Mylene is 20 years older than her daughter Hazel. In 10 years time she will be twice as older as her daughter. How older are Mylene and Hazel?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52786)

.
Mylene is 20 years older than her daughter Hazel.
In 10 years time she will be twice as older as her daughter. How older are Mylene and Hazel?
~~~~~~~~~~~~~~

If H is x years old, then, according to the condition, M is (x+20) years old.


In 10 years


    (x+20) + 10 = 2*(x + 10).


Simplify and find x


    x + 30 = 2x + 20


    30 - 20 = 2x - x


      10    =   x


ANSWER.  H is 10 years old.   M is 10+20 = 30 years old.


CHECK.  In 10 years,  H will be 20 years old;  M will be 40 years old.

        In 10 years, M will be twice the age of H.   ! Correct !

Solved.

----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
    - Entertainment age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.




Answer by Alan3354(69443)

Mylene is 20 years older than her daughter Hazel. In 10 years time she will be twice as older as her daughter. How older are Mylene and Hazel?
---
"twice as older" makes no sense.
============
Assuming you meant twice as old:
--------------
The age difference is constant at 20 years.
---> when they're 20 & 40.